OpenStudy (anonymous):

solve the system of equations using elimination 2x+4y=-8 -3x+4y=2

OpenStudy (anonymous):

change the sign on one of the equations and you can get rid of the y

OpenStudy (theviper):

Subtract second equation from first:- 2x+4y=-8 -(-3x)-(4y)=-2 ------------- 5x=-10 x=-10/5=2

OpenStudy (theviper):

Now put the value of 'x' in any equation & find 'y' :)

OpenStudy (theviper):

Can u do that from now???

OpenStudy (anonymous):

y=-16 ?

OpenStudy (anonymous):

and x=-2 ?

OpenStudy (theviper):

Wait I have done a mistake it's x=-2 2x+4y=-8

OpenStudy (theviper):

2(-2)+4y=-8 -4+4y=-8 4y=-8+4 4y=-4 y=-4/4=-1

OpenStudy (anonymous):

so y=-1 ?

OpenStudy (theviper):

yes why not ;)

OpenStudy (theviper):

Understand now :)

OpenStudy (anonymous):

sorry it took so long to reply, i got busy but i got x=2, y=3

OpenStudy (theviper):

See my solution @jkd13 :)

OpenStudy (anonymous):

haha we made the same mistake but yes your right.

OpenStudy (anonymous):

your second answer

OpenStudy (theviper):

it's right:)